Miami Children's Hospital reported net income for 2018 0f $9,700,120 on total revenues of $121,251,500. Depreciation expense totaled $4,041,717. Assume that all expenses, except depreciation, were cash expenses.

a. Assume the hospital is a not-for-profit corporation for parts a - c. What were total operating expenses for 2018?

b. What were total cash expenses for 2018?

c. What was the hospital’s 2018 cash flow?

d. Now, suppose the company changed its depreciation method (still within GAAP) such that its depreciation expense double. How would this change affect cash flow? Please show your work.

e. Assume the hospital is a for-profit corporation who reported a pretax (operating) income of $9,700,120 on total revenue of $121,251,500 and that depreciation totaled $4,041,717, and that all expenses, except depreciation, were cash expenses. What was the hospital’s 2018 cash flow if the tax rate is 20% of pretax income?

f. Assume the hospital is a for-profit corporation as in Part e above. Suppose the company changed its depreciation method (still within GAAP) such that its depreciation expense double. How would this change affect cash flow? Please show your work.
Answer #1

A. Total operating expenses = Revenue - net income = 121,251,500-9,700,120 = 111,551,380

B. Total Cash expenses = operating expenses - depreciation = 111,551,380 - 4,041,717 = 107,509,663

C. Cash flow from operating = net income + non cash expenses + changes in working capital

CF = 9,700,120 + 4,041,717 = 13,741,837

D. New depreciation = 8,083,434

CF = 9700120+8083434= 17,783,554

E. Net income = 9700120*(1-0.2) = 7,760,096

Cash flow = 7760096+4041717=11,801,813

F. New depreciation = 8083434

CF = 7760096+8083434= 15,843,530